# Laminar chaos in systems with random and chaotically time-varying delay

### v1

@@ I Introduction
where the variable delay $\tau(\bm{h}(t))$ depends on the output $\bm{h}(t)$ of an independent delay generating process. For constant delay, $\tau(\bm{h}(t))=\tau_{0}$, such systems were extensively studied, where the choice of nonlinearity $f$ of the feedback depends on the specific application. The nonlinearity $f(z)=\mu\,z/(1+z^{10})$ characterizes the Mackey-Glass equation [62] which is a blood production model, a sinusoidal nonlinearity, $f(z)=\mu\,\sin(z)$, gives the Ikeda equation [63, 64], which is a model for light dynamics in a ring cavity with a nonlinear optical medium and describes certain optoelectronic oscillators [49, 65, 66]. General properties can be derived using a simpler quadratic nonlinearity $f(z)=\mu\,z(1-z)$ [67]. Normal and anomalous diffusion was found in systems with the climbing-sine nonlinearity $f(z)=z+\mu\,\sin(2\pi\,z)$ [68, 69] and the double-sine nonlinearity $f(z)=z-\mu[\sin(2\pi\,z)+(1/2)\sin(4\pi\,z)]$ [70], respectively. The parameter $\Theta$ sets the overall timescale of the system. If $\Theta$ is large such that the short time scale $(1/\Theta)$ of the system is much smaller than the delay, Eq. (1) belongs to the class of singularly perturbed delay differential equations and large delay systems, which are widely studied for constant delay, where slowly oscillating periodic solutions [71, 72, 73, 67], high-dimensional chaotic dynamics [74, 75], chaos control [76], spatio-temporal phenomena [77] and chaotic diffusion [69] have been observed. Also results on systems with state-dependent delay are available [78, 79, 80, 81, 82]. The potentially high-dimensional dynamics of systems with the structure of Eq. (1) and similar singularly perturbed systems is interesting for applications such as chaos communication [83, 84, 85, 86], random number generation [87, 88, 89], and reservoir computing [90, 91, 92, 93].
@@
The paper is structured as follows. In Sec. II we provide the definition of the considered random and chaotic delay variations. A short review on the theory of laminar chaos and the concept of conservative and dissipative delays is given in Sec. III. The generalization of the concept of laminar chaos is provided in Sec. IV, where the dynamics of the delay system with respect to the delay parameter space is first analyzed numerically. The results are then explained and verified by analytical considerations, which are exact in the large delay limit $\Theta\to\infty$, using the theory of circle maps with quenched disorder provided in [94]. In Sec. V characteristic features of laminar chaos in systems with random and chaotically time-varying delay are identified and the test for laminar chaos introduced in [50, 95] is generalized and benchmarked.
@@ II Random and chaotically time-varying delay
In this section, we define the notion of random and chaotically time-varying delays that will be explored in this manuscript. The original theory on laminar chaos was derived for periodic delay variations, where a sinusoidal delay of the form,
@@
served as a generic example. This parameterization with the mean delay $\tau_{0}$ and the amplitude $A$ when used in Eq. (1) enables a systematic analysis of observables in the delay parameter space formed by $A$ and $\tau_{0}$, revealing interesting features such as a huge variation of the attractor dimension as a function of $\tau_{0}$ in a fractal manner [47]. We replace the sinusoidal delay with random and chaotically varying functions $\tau(t)=\tau_{\omega}(t)$, while keeping this form of parameterization. The index $\omega$ represents the disorder parameter and therefore determines the specific realization of the random delay or the set of initial conditions of the chaotic delay generating process. As done in Eq. (2) the delay variation is normalized such that the amplitude $A$ can be varied between 00 and $1$, where it is ensured that we always have
@@
In principle, a random delay can be defined as the limit of a quasiperiodic delay with an infinite number of randomly chosen frequencies leading to first interesting results [54]. However, to fulfill Eq. (3) in the generic case of a sum of cosines with random frequencies as considered in [54], the overall amplitude of the delay tends to zero as the number of frequencies increases such that the limit of an infinite number of frequencies leads to a constant delay. Therefore, in this paper, we use a different approach to generate random time-varying delays. For our numerical analysis, we consider three examples of generic smooth random functions, which are illustrated in Fig. 1. The differences between them allow us to distinguish between general results and more specific results that follow from the structure of the delay variation. We choose our random delays $\tau_{\omega}(t)$ to be a smoothed, scaled and shifted version of a piecewise linear signal $\chi_{\omega}(t)$, which determines the statistical properties of the delay variation. We choose three different $\chi_{\omega}(t)$ signals. In Fig. 1(a), we first choose $\chi_{\omega}(t)$ is a piecewise constant dichotomic signal with values $S_{i}=\pm 1$ and a residence time equal to one, i.e., we have
@@
where $i$ are integers. For the second example, as shown in Fig. 1(b), $\chi_{\omega}(t)$ is given by a periodic sawtooth wave, which is randomly switched on and off. We define it for $t\in[i-1,i)$ as
@@
where the constant $c$ is chosen such that the time average of $\chi_{\omega}(t)$ vanishes for almost all realizations. In both cases a specific realization of $\chi_{\omega}(t)$ is determined by a symbol chain $\omega=S_{1}S_{2}\ldots$, given by an infinite number of random symbols $S_{i}$, which are independent and distributed with equal probability. The third example shown in Fig. 1 is inspired by the sawtooth-shaped random delay in [57]. In this case, $\chi_{\omega}(t)$ is given by a sawtooth wave that consists of piecewise linear functions with slope equal to one, where the $i$th linear segment lasts for a randomly chosen time $S_{i}$ and then, after resetting to a fixed initial value, the next linear segment with duration $S_{i+1}$ begins. We have for $t\in[T_{i-1},T_{i})$
@@
and $T_{0}=0$, where the constant $c$ is chosen such that the time average of $\chi_{\omega}(t)$ vanishes for almost all realizations. The $S_{i}$ are independent and uniformly distributed in $[1/2,3/2]$, where the mean duration of the linear segments equals the duration of the linear segments in the on-off sawtooth delay. In all cases, the random delay variation is then obtained by a convolution $(G_{\varsigma}\ast\chi_{\omega})(t)$ with a Gaussian $G_{\varsigma}(t)$ of variance $\varsigma^{2}$. Scaling by $AA_{s}$ and shifting by an offset $\tau_{0}$ gives our random delay
@@
[…]
@@ III Review of laminar chaos
[…]
which is the (nonlinear) solution operator that generates the solution segment $z_{k+1}(t)$ from the preceding solution segment $z_{k}(t)$. The low pass filter is now expressed as a smoothing operator with the exponential kernel $\Theta e^{-\Theta(t-t^{\prime})}$ with width $(1/\Theta)$. In this work we are interested in the limit of a large cutoff frequency $\Theta$ and therefore we consider $\Theta\to\infty$, where the kernel converges to a delta distribution so that the smoothing operator becomes the identity [75]. We then obtain the so-called limit map given by
@@
which is also a good approximation of Eq. (11) for finite $\Theta$ given that $\dot{z}(t)$ is much smaller than $\Theta$. In terms of the smoothing operator this means that the approximation is accurate if the timescale of the solution $z(t)$ is much larger than the width $(1/\Theta)$ of the kernel. The dynamics of the solution segment $z_{k}(t)$ governed by the limit map, Eq. (12), can be interpreted as iteration of the graph $(t,z_{k}(t))$ under the two dimensional map given by
@@
which consists of two independent one-dimensional maps. Equation (13a) is the inverse of the access map given by Eq. (10) and Eq. (13b) is a map defined by the nonlinearity $f$ of the feedback. As it was found for general delay systems with one periodic or quasiperiodic delay [45, 46, 54], the influence of the time-varying delay is governed by the so-called access map defined by Eq. (10). For periodic or quasiperiodic delay, the access map is the lift of a circle map (cf. [101]) or a foliation preserving torus map (cf. [102, 103]), respectively. Based on the dynamics of these maps there are two universal classes of time-varying delays. If the access map shows marginally stable dynamics such that its Lyapunov exponent vanishes, $\lambda[R]=0$, and perturbations of a reference orbit do on average neither grow nor decay, the delay belongs to the class of conservative delays. Such delays are equivalent to constant delays in the sense that the delay system can be transformed into a system with constant delay but time-dependent coefficients. This means that they show qualitatively the same dynamics. In contrast, so-called dissipative delays belong to stable access map dynamics, where the Lyapunov exponent of the access map is negative, $\lambda[R]<0$. Such delays are not equivalent to constant delay systems, and thus they can lead to drastic differences in the dynamics compared to constant delay systems.
@@
These differences become very pronounced if one considers chaotic dynamics of Eq. (1) in the limit of large $\Theta$. Chaos generated by this systems is weak chaos in the sense of [104] (not to be confused with weak chaos from chaotic diffusion [70]), which originates from the delayed feedback due to the negative instantaneous Lyapunov exponent. Given that the map defined by the feedback nonlinearity, Eq. (13b), is chaotic, stretching and folding of the function values of the signal leads to strong fluctuations as the signal circles inside the feedback loop. If the delay is a conservative delay, it only leads to an additional frequency modulation due to the Doppler effect induced by the delay variation, which can not compensate the stretching an folding induced by the feedback nonlinearity. Therefore the characteristic frequency of the signal increases with each roundtrip inside the feedback loop and is only bounded by the cut-off frequency $\Theta$ of the low-pass filter. This basically is the mechanism behind so-called turbulent chaos [75], which is also found in constant delay systems and is named after the optical turbulence found for an optical ring cavity in [64]. It is a very high-dimensional dynamics as the dimension of the chaotic attractor is proportional to $\Theta$ and thus can be arbitrarily large [74]. In stark contrast, dissipative delays can lead to comparably low-dimensional chaos as demonstrated in [47, 48, 54]. The main mechanism behind this is the so-called resonant Doppler effect, where the frequency modulation due to the delay variation is in resonance with the average roundtrip time inside the feedback loop - a consequence of the mode-locking behavior of the access map. The competition between the resonant Doppler effect and the stretching and folding of the feedback nonlinearity leads to the development of low frequency phases that are periodically or quasiperiodically interrupted by rather short high frequency phases, which characterize the comparably low-dimensional types of chaos called (generalized) laminar chaos. If the condition for laminar chaos
$\lambda[f]+\lambda[R]<0$ (14)
@@
is fulfilled, where $\lambda[f]$ is the Lyapunov exponent of the map defined by Eq. (13b), the low frequency phases degenerate to nearly constant laminar phases, given that their duration is much larger than the width $(1/\Theta)$ of the smoothing kernel in Eq. (11). Since the derivative of these laminar phases is small, the two-dimensional map, Eq. (13) gives a good approximation of the dynamics of these phases, where (13a) and (13b) governs the dynamics of the durations and the levels of the laminar phases, respectively. Neglecting the high-frequency phases, which degenerate to short burstlike transitions between the laminar phases, laminar chaos can be described by this low-dimensional system leading to a very low-dimensional chaotic attractor. If $\lambda[R]$ is negative but Eq. (14) is not fulfilled, generalized laminar chaos of order $(m-1)$ is observed, where $m>0$ is the smallest integer such that
@@
holds. In this case, $(m-1)$-order polynomials take the role of the nearly constant phases of classical laminar chaos leading to an effective dimension of order $O(m)$, which is small if $m$ is small [48].
In this paper, we generalize this theory to systems with random and chaotically time-varying delay. Using the theory of disordered circle maps provided in [94], we demonstrate that laminar chaos can be observed in such systems. Further we show that the simple condition for laminar chaos, Eq. (14) is still valid. This is possible since access maps for random delay can be treated as infinite size limit, $L\to\infty$, of disordered circle maps with size and period $L$ given by
@@
where the delay $\tau(t)$ is defined as a random function in an interval of length $L$, which is then repeated periodically. So for all finite sizes $L$, we can rely on the known properties of periodic delays, while approaching the infinite size limit $L\to\infty$. In this limit, almost all realizations of the random delay lead to the same well-defined Lyapunov exponent $\lambda[R]$. Therefore, Eqs. (14,15) are directly applicable since they were derived directly from a stability analysis of polynomial solutions of the limit map, Eq. (12), only assuming that the maps defined by $f$ and $R$ have a well defined Lyapunov exponent (see [47, 48] for details). While the randomness of the delay variation has no influence on the validity of the criteria for (generalized) laminar chaos, it drastically changes the delay parameter space in the sense that these low-dimensional types of chaos are observed in almost the whole parameter space, such that high-dimensional turbulent chaos becomes extremely rare. In Sec. V we also find interesting changes in the dynamical properties of laminar chaos compared to periodic and quasiperiodic delay systems, which make it necessary to update the time series analysis toolbox for laminar chaos developed in [49, 50, 95].

### v2

@@ I Introduction
where the variable delay $\tau(\bm{h}(t))$ depends on the output $\bm{h}(t)$ of an independent delay generating process. Systems with constant delay, $\tau(\bm{h}(t))=\tau_{0}$, have been extensively studied, where the choice of nonlinearity $f$ of the feedback depends on the specific application. The nonlinearity $f(z)=\mu\,z/(1+z^{10})$ characterizes the Mackey-Glass equation [62] which is a blood production model. A sinusoidal nonlinearity, $f(z)=\mu\,\sin(z)$, gives the Ikeda equation [63, 64], which is a model for light dynamics in a ring cavity with a nonlinear optical medium and describes certain optoelectronic oscillators [49, 65, 66]. General properties can be derived using a simpler quadratic nonlinearity $f(z)=\mu\,z(1-z)$ [67]. Normal and anomalous diffusion were found in systems with the climbing-sine nonlinearity $f(z)=z+\mu\,\sin(2\pi\,z)$ [68, 69] and the double-sine nonlinearity $f(z)=z-\mu[\sin(2\pi\,z)+(1/2)\sin(4\pi\,z)]$ [70], respectively. The parameter $\Theta$ sets the overall timescale of the system. If $\Theta$ is large such that the short time scale $(1/\Theta)$ of the system is much smaller than the delay, Eq. (1) belongs to the class of singularly perturbed delay differential equations and large delay systems, which are widely studied for constant delay, where slowly oscillating periodic solutions [71, 72, 73, 67], high-dimensional chaotic dynamics [74, 75], chaos control [76], spatio-temporal phenomena [77] and chaotic diffusion [69] have been observed. Also results on systems with state-dependent delay are available [78, 79, 80, 81, 82]. The potentially high-dimensional dynamics of systems with the structure of Eq. (1) and similar singularly perturbed systems is interesting for applications such as chaos communication [83, 84, 85, 86], random number generation [87, 88, 89], and reservoir computing [90, 91, 92, 93].
@@
In this paper, we generalize the concept of laminar chaos to systems with random and chaotically time-varying delays. The generic delay variations used for numerics are defined in Sec. II. After reviewing the theory on laminar chaos in Sec. III, including the concept of conservative and dissipative delays for (quasi-)periodic delays, their generalizations are provided in Sec. IV. We first numerically analyze the dynamics for short-time correlated random and chaotically time-varying delays, where we present our main finding: Low-dimensional laminar chaos and its generalizations can be found in almost the whole delay parameter space spanned by mean and amplitude of the delay, whereas high-dimensional turbulent chaos is suppressed. The results are then explained and verified by analytical considerations, which are exact in the large delay limit $\Theta\to\infty$, using the theory of circle maps with quenched disorder provided in [94]. In particular, we find that the conditions for (generalized) laminar chaos presented in Sec. III are still valid for random and chaotically time-varying delays because relevant physical quantities are unique for almost all realizations. Our main finding is confirmed by the fact that these conditions are fulfilled if the minimum of the delay exceeds the correlation length of the delay variation leading to (generalized) laminar chaos above that threshold, i.e., almost everywhere. In Sec. V characteristic features of laminar chaos in systems with random and chaotically time-varying delay are identified and the test for laminar chaos introduced in [50, 95] is generalized and benchmarked.
@@ II Random and chaotically time-varying delay
In this section, we define the notion of random and chaotically time-varying delays that will be explored in this manuscript. The original theory on laminar chaos was derived for periodic delay variations, where a sinusoidal delay of the form
@@
served as a generic example. This parameterization with the mean delay $\tau_{0}$ and the amplitude $A$ when used in Eq. (1) enables a systematic analysis of observables in the delay parameter space formed by $A$ and $\tau_{0}$, revealing interesting features such as a huge variation of the attractor dimension as a function of $\tau_{0}$ in a fractal manner [47]. We replace the sinusoidal delay with random and chaotically varying functions $\tau(t)=\tau_{\omega}(t)$, while keeping this form of parameterization. The index $\omega$ represents the disorder parameter and therefore determines the specific realization of the random delay or the set of initial conditions of the chaotic delay generating process. As done in Eq. (2) the delay variation is normalized such that the amplitude $A$ can be varied between 0 and $1$, where it is ensured that we always have
@@
For our numerical analysis, we consider three examples of generic smooth random functions, which are illustrated in Fig. 1. The differences between them allow us to distinguish between general results and more specific results that follow from the structure of the delay variation. Each random delay $\tau_{\omega}(t)$ is smoothed, scaled and shifted version of a piecewise linear signal $\chi_{\omega}(t)$, which determines the statistics  [^1] In principle, a random delay can be defined as the limit of a quasiperiodic delay with an infinite number of randomly chosen frequencies leading to first interesting results [54]. However, to fulfill Eq. (3) in the generic case of a sum of cosines with random frequencies as considered in [54], the overall amplitude of the delay tends to zero as the number of frequencies increases such that the limit of an infinite number of frequencies leads to a constant delay. We therefore use a different approach.. We choose three different signals $\chi_{\omega}(t)$:
Dichotomic delay
$\chi_{\omega}(t)$ is a piecewise constant dichotomic signal with values $S_{i}=\pm 1$ and a residence time equal to one, i.e., we have
@@
where $i$ are integers.
On-off sawtooth delay
$\chi_{\omega}(t)$ is given by a periodic sawtooth wave, which is randomly switched on and off. We define it for $t\in[i-1,i)$ as
@@
where the constant $c$ is chosen such that the time average of $\chi_{\omega}(t)$ vanishes for almost all realizations.
In cases (a) and (b), a realization is determined by an infinite symbol chain $\omega=S_{1}S_{2}\ldots$ of independent random symbols $S_{i}$, which are distributed with equal probability.
Random-duration sawtooth
Inspired by the sawtooth-shaped random delay in [57], $\chi_{\omega}(t)$ is given by a sawtooth wave with slope equal to one, where the $i$th linear segment lasts for a randomly chosen time $S_{i}$ and then, after resetting to a fixed initial value, the next linear segment with duration $S_{i+1}$ begins. We have for $t\in[T_{i-1},T_{i})$
@@
and $T_{0}=0$, where the constant $c$ is chosen such that the time average of $\chi_{\omega}(t)$ vanishes for almost all realizations. The $S_{i}$ are independent and uniformly distributed in $[1/2,3/2]$, where the mean duration of the linear segments equals the duration of the linear segments in the on-off sawtooth delay.
In all cases, the random delay variation is then obtained by a convolution $(G_{\varsigma}\ast\chi_{\omega})(t)$ with a Gaussian $G_{\varsigma}(t)$ of variance $\varsigma^{2}$. Scaling by $AA_{s}$ and shifting by an offset $\tau_{0}$ gives our random delay
@@
The first three equations make up the Lorenz system [99], where we use the parameters $\sigma=10$, $\rho=28$, $\beta=\frac{8}{3}$ for chaos as used originally by Lorenz. The constant $\bar{h}_{4}=\lim_{T\to\infty}\frac{1}{T}\int_{t_{0}}^{t_{0}+T}dt\,h_{4}(t)$ ensures that the time average vanishes. The delay variation is given by the solution of Eq. (9d), which is a filtered version of the derivative $\dot{h}_{1}$ of the first component of the Lorenz system. The nonlinear filter is designed such that the delay variation mimics the random duration sawtooth delay shown in Fig. 1(c)  [^2] The sawtooth shape appears since we have $\dot{h}_{4}\approx 1$ for $\gamma\dot{h}_{1}\gg 0$ and $\dot{h}_{4}\approx 1+\gamma\,\dot{h}_{1}\,h_{4}$ for $\gamma\dot{h}_{1}\ll 0$. This means that $h_{4}(t)$ increases approximately linearly with slope $1$ if $\gamma\dot{h}_{1}\gg 0$ and rapidly decreases if $\gamma\dot{h}_{1}\ll 0$., since for sawtooth-shaped variations the strongest effect of the time-varying delay is observed as we will in Sec. IV.1 and Sec. IV.2. However, we note that the theory presented here is not restricted to such variations and is valid for all short time correlated chaotically varying delays. In the following numerical analysis we choose the filter parameter $\gamma=0.2$, which leads to the chaotically time-varying delay shown in Fig. 2(b), where the unfiltered chaotic signal $\dot{h}_{1}(t)$ is shown in Fig. 2(a).
@@ III Review of laminar chaos
$z_{k+1}(t)=z_{k}(t_{k})e^{-\Theta(t-t_{k})}+\int\limits_{t_{k}}^{t}\!dt^{\prime}\,\Theta e^{-\Theta(t-t^{\prime})}f(z_{k}\bm{(}R(t^{\prime})\bm{)}),$ (11)
a (nonlinear) solution operator generating the solution segment $z_{k+1}(t)$ from the preceding solution segment $z_{k}(t)$. The low pass filter acts as a smoothing operator with exponential kernel $\Theta e^{-\Theta(t-t^{\prime})}$ of width $(1/\Theta)$. We focus on the limit of a large cutoff frequency $\Theta$ and therefore we consider $\Theta\to\infty$, where the kernel converges to a delta distribution so that the smoothing operator becomes the identity [75]. We then obtain the so-called limit map given by
@@
which is also a good approximation of Eq. (11) for finite $\Theta$ given that $\dot{z}(t)\ll\Theta$, i.e, when the timescale of the solution $z(t)$ is much larger than the width $(1/\Theta)$ of the kernel. Under the limit map, the evolution of the segment $z_{k}(t)$ can be interpreted as iteration of the graph $(t,z_{k}(t))$ by the two dimensional map
@@
which consists of two independent one-dimensional maps. Equation (13a) is the inverse of the access map given by Eq. (10) and Eq. (13b) is a map defined by the nonlinearity $f$ of the feedback. For periodic or quasiperiodic delay, the access map is the lift of a circle map (cf. [103]) or a foliation preserving torus map (cf. [104, 105]), respectively. For general systems with (quasi-)periodic delay it is known that the dynamics of these maps define two universal classes of time-varying delays [45, 46, 54]. If the dynamics is marginally stable dynamics with vanishing Lyapunov exponent, $\lambda[R]=0$, and perturbations of a reference orbit do on average neither grow nor decay, the delay belongs to the class of conservative delays. Such delays are equivalent to constant delays in the sense that the delay system can be transformed into a system with constant delay but time-dependent coefficients. This means that they show qualitatively the same dynamics. In contrast, dissipative delays belong to stable access map dynamics, $\lambda[R]<0$. Such delays are not equivalent to constant delay systems, and thus they can lead to drastic differences in the dynamics compared to constant delay systems.
@@
These differences become very pronounced for chaotic dynamics of Eq. (1) at large $\Theta$. Chaos here is weak in the sense of [106] (not to be confused with weak chaos from chaotic diffusion [70]) as it originates from the delayed feedback due to the negative instantaneous Lyapunov exponent. Given that the map defined by the feedback nonlinearity, Eq. (13b), is chaotic, stretching and folding of the function values of the signal leads to strong fluctuations as the signal circles inside the feedback loop. For conservative delays, the delay variation induces only an additional frequency modulation due to the Doppler effect, which can not compensate the stretching and folding induced by the feedback nonlinearity. Therefore the characteristic frequency of the signal increases with each roundtrip inside the feedback loop and is only bounded by the cut-off frequency $\Theta$ of the low-pass filter. This is the mechanism behind turbulent chaos [75], also found in constant delay systems and named after optical turbulence in an optical ring cavity [64]. It is a very high-dimensional dynamics as the dimension of the chaotic attractor is proportional to $\Theta$ and thus can be arbitrarily large [74]. In stark contrast, dissipative delays can lead to comparably low-dimensional chaos as demonstrated in [47, 48, 54]. The main mechanism behind this is the so-called resonant Doppler effect, where the frequency modulation due to the delay variation is in resonance with the average roundtrip time inside the feedback loop - a consequence of the mode-locking behavior of the access map. The competition between the resonant Doppler effect and the stretching and folding of the feedback nonlinearity leads to the development of low frequency phases that are periodically or quasiperiodically interrupted by rather short high frequency phases, which characterize the comparably low-dimensional types of chaos called (generalized) laminar chaos. Laminar chaos occurs when
$\lambda[f]+\lambda[R]<0,$ (14)
@@
where $\lambda[f]$ is the Lyapunov exponent of the map defined by Eq. (13b). Then the low frequency phases degenerate to nearly constant laminar phases provided their duration is much larger than the width $(1/\Theta)$ of the smoothing kernel in Eq. (11). Since the derivative of these laminar phases is small, the two-dimensional map, Eq. (13) well approximates their dynamics, where (13a) and (13b) governs the dynamics of the durations and the levels of the laminar phases, respectively. Neglecting the high-frequency phases, which degenerate to short burstlike transitions between the laminar phases, laminar chaos can be described by this low-dimensional system leading to a very low-dimensional chaotic attractor. If $\lambda[R]$ is negative but Eq. (14) is not fulfilled, generalized laminar chaos of order (m1)(m-1) is observed, where $m>0$ is the smallest integer such that
@@
holds. In this case, (m1)(m-1)-order polynomials take the role of the nearly constant phases of classical laminar chaos leading to an effective dimension of order $O(m)$, which is small if $m$ is small [48].
In this paper, we generalize this theory to systems with random and chaotically time-varying delay. Using the theory of disordered circle maps provided in [94], we demonstrate that laminar chaos can be observed in such systems and the condition for laminar chaos, Eq. (14), remains valid. We use that access maps for random delay can be treated as infinite size limit, $L\to\infty$, of disordered circle maps with size and period $L$ given by
@@
where the delay $\tau(t)$ is defined as a random function in an interval of length $L$ and repeated periodically. So for all finite sizes $L$, we can rely on the known properties of periodic delays, while approaching the infinite size limit $L\to\infty$. In this limit, almost all realizations of the random delay lead to the same well-defined Lyapunov exponent $\lambda[R]$. Hence Eqs. (14,15) apply directly, as they follow from a stability analysis of polynomial solutions of the limit map, Eq. (12), only assuming that the maps defined by $f$ and $R$ have a well defined Lyapunov exponent (see [47, 48] for details). While randomness of the delay variation does not affect the validity of the criteria for (generalized) laminar chaos, it drastically changes the delay parameter space: These low-dimensional types of chaos occupy almost all of it, such that high-dimensional turbulent chaos becomes extremely rare. In Sec. V we also find interesting changes in the dynamical properties of laminar chaos compared to periodic and quasiperiodic delay systems, which make it necessary to update the time series analysis toolbox for laminar chaos developed in [49, 50, 95].